OpenStudy (anonymous):

An automobile dealer had 250 vehicles on his lot during the month of June. He must pay a monthly inventory tax of $3 per car and $4 per truck. If his tax bill in June was $850, then find the number of cars and the number of trucks he had on his lot in June.

OpenStudy (texaschic101):

c + t = 250 ----> c = 250 - t now sub 250 - t in for c in the other equation 3c + 4t = 850 3(250 - t) + 4t = 850 -- distribute 750 - 3t + 4t = 850 -- subtract 750 -3t + 4t = 850 - 750 -- combine like terms t = 100 now sub 100 in for t in either equation to find c c + t = 250 c + 100 = 250 c = 250 - 100 c = 150 check.. 3c + 4t = 850 3(150) + 4(100) = 850 450 + 400 = 850 850 = 850 (correct) there are 150 cars (c) and there are 100 trucks (t) any questions ?

OpenStudy (anonymous):

i think that pretty much explains my question thank you i always have trouble with setting my problem up but this was really easy to understand